Problem 8-8 Morgans Televisions produces television sets in three categories: portable, midsize, and flat-screen. On January 1, 2017, Morgan adopted dollar-value LIFO and decided to une a single inventory pool. The companys January 1 inventory consists of: Category Quantity Cost per Unit Total Cost Portable Midsize Flat-screen 5,700 8,100 3.100 16,900 115 655,500 288 2,332,800 460 1,426,000 $4,414,300 During 2017, the company had the following purchases and sales Quantity Quantity Selling Price Category Purchased Cost per Unit Sold Portable Midsize per Unit 15,200 20,900 9,800 $127 13,500 345 24,600 75 5,700 3,800 $173 466 690 Flat-screen90 45,900
Your answer is incorrect. Try again. Compute ending inventory, cost of goods sold, and gross profit. (Round answers to o decimal places, e.g. 6,548.) Ending inventory Cost of goods sold Gross profit LINK TO TEXT
Your answer is incorrect. Try again. Assume the company uses three inventory pools instead of one. Compute ending inventory, cost of goods sold, and gross profit.(Round price index to 6 decimal places e.g. 1.456287 and final answers to O decimal places, e.g. 6,548.1 Ending inventory Cost of goods sold Gross profit LINK TO TEXT
Answer #1
  1. Ending Inventory=$ 5,922,100( Refer to table and note below)
  2. Cost of Goods Sold=$ 13,479,000( Refer to table and note below)
  3. Gross Profit=$ 4,253,100( Refer to table and note below)

Particulars

Portable

Midsize

Flat -screen

Opening stock

5700

8100

3100

Add: Purchases

15200

20900

9800

20900

29000

12900

Less: Goods Sold

13500

24600

5700

Closing Stock

7400

4400

7200

Cost of closing stock

1700 units*$127

5700 units*$115

4400units*288

4100units*575

3100units*460

Cost of closing stock

$(215900+655500)

=$ 8,71,400

$ 1,267,200

$(2357500+1426000)

$ 3,783,500

Selling Price per unit

$ 173

$ 466

$ 690

Costof goods sold/unit

$ 127

$ 345

$ 575

Gross Profit per unit

$ 46

$ 121

$115

Total Cost of goods sold

$127*13500

=$ 1,714,500

$ 345*24600

=$ 8,487,000

$575*5700

=$3,277,500

Total Gross Profit

$46*13500

=$6,21,000

$121*24600

=$ 2,976,600

$115*5700

=$ 6,55,500

Note :

  1. Ending Inventory=$(871400+1267200+3783500)=$ 5,922,100
  2. Cost of Goods Sold=$(1714500+8487000+3277500)=$13,479,000
  3. Gross Profit=$(621000+2976600+655500)=$4,253,100